### v1

# J-CHAT: Japanese Large-scale Spoken Dialogue Corpus for Spoken Dialogue Language Modeling

Abstract
Spoken dialogue plays a crucial role in human-AI interactions, necessitating dialogue-oriented spoken language models (SLMs). To develop versatile SLMs, large-scale and diverse speech datasets are essential. Additionally, to ensure hiqh-quality speech generation, the data must be spontaneous like in-wild data and must be acoustically clean with noise removed. Despite the critical need, no open-source corpus meeting all these criteria has been available. This study addresses this gap by constructing and releasing a large-scale spoken dialogue corpus, named Japanese Corpus for Human-AI Talks (J-CHAT), which is publicly accessible. Furthermore, this paper presents a language-independent method for corpus construction and describes experiments on dialogue generation using SLMs trained on J-CHAT. Experimental results indicate that the collected data from multiple domains by our method improve the naturalness and meaningfulness of dialogue generation.
J-CHAT: Japanese Large-scale Spoken Dialogue Corpus
for Spoken Dialogue Language Modeling
Wataru Nakata1∗  Kentaro Seki1 thanks:  Equal contribution.  Hitomi Yanaka1 Yuki Saito1Shinnosuke Takamichi1,2Hiroshi Saruwatari1 1 The University of Tokyo, Japan  2 Keio University, Japan nakata-wataru855@g.ecc.u-tokyo.ac.jp
1 Introduction
Spoken dialogue systems have garnered significant attention for their role in enabling human-AI communication Mitsui et al. (2023). Spoken language holds rich expressiveness through tone, intonation, and non-verbal vocalizations such as laughter and screams, highlighting the importance of spoken dialogue systems. To leverage the expressive capabilities of spoken language in spoken dialogue systems, it is essential to develop spoken language models (SLMs) that model the linguistic structures of spoken dialogue. In other words, SLMs are crucial for facilitating smooth and effective human-AI communication. Generative spoken language modeling (GSLM, Lakhotia et al., 2021) using HuBERT Hsu et al. (2021) as a speech feature extractor is a well-known example of SLMs.
Training SLMs requires a large-scale speech corpus. Open-source corpora are crucial because they lower research barriers and ensure reproducible results. Despite their importance, the development of spoken dialogue corpora has been limited. Current methods for constructing spoken dialogue corpora Cieri et al. (2004); Saito et al. (2022) are labor-intensive making it challenging to scale the corpora. Furthermore, large-scale spoken dialogue corpora have primarily been constructed for English Cieri et al. (2004), with limited progress in other languages, such as Japanese.
Therefore, we construct and release a new open-source corpus named “Japanese Corpus for Human-AI Talks” (J-CHAT) [^1] https://huggingface.co/datasets/sarulab-speech/J-CHAT. To collect large-scale speech data, we propose a fully automated thus scalable method for corpus construction, including data collection from the web. Gathering in-the-wild data ensures the collection of spontaneous speech covering a wide range of conversational topics. Additionally, we use noise removal techniques to make the data acoustically clean. Our contributions include:
  • Constructing and releasing a novel, open-source, and large-scale speech corpus.
  • Providing a spontaneous and clean speech corpus, thereby facilitating the generation of natural dialogues.
  • Developing a corpus construction method that is independent of the language.
Table 1: Comparison of speech corpora related to this study, sorted by size.
Corpus name Size(hours) Open-source Dialogue Spontaneous Clean speech
STUDIES (Saito et al., 2022) $8.2$
& ✅   ✅ DailyTalk (Lee et al., 2023) $20$ Fisher (Cieri et al., 2004) $2k$   ✅ ✅ ✅ GigaSpeech (Chen et al., 2021) $33k$   -   J-CHAT (This study) $69k$
2 Related Work
2.1 Spoken Language Models
Large-scale SLMs have been reported to enhance the performance of various downstream tasks. Training these models requires corpora that are significantly larger than traditional ones. For instance, research using 100k hours of audio for dialogue generation Borsos et al. (2023), 680k hours for speech recognition Radford et al. (2023), and 4.5M hours for speech translation Barrault et al. (2023) has been reported. These studies demonstrate that increasing the amount of data leads to performance improvements, motivating us to construct large-scale speech corpora.
However, these studies primarily focus on model architecture improvements, and the datasets used in the experiments are often not publicly available, with little to no detail on their construction methods. One research problem that suffers from a lack of data is dialogue spoken language modeling Nguyen et al. (2023). In a previous study, 2000 hours of dialogue data were used, drawing from one of the largest available dialogue speech data resources. Despite this substantial dataset, the researchers reported that its size was insufficient for effective dialogue spoken language modeling. By proposing and releasing a method for constructing a large-scale corpus specifically for the dialogue domain, we aim to facilitate discussions and advancements in corpus construction techniques.
2.2 Existing Speech Corpora
Training SLMs for dialogues requires large-scale, spontaneous and acoustically clean dialogue speech data. Table 1 shows corpora related to this study. STUDIES corpus Saito et al. (2022) and DailyTalk corpus Lee et al. (2023) are open-source dialogue corpora designed for multi-turn conversations, but they are very small in size because they are manually recorded spoken dialogues. Fisher corpus Cieri et al. (2004) is more extensive but is not open-source. GigaSpeech corpus Chen et al. (2021) is a large-scale open-source corpus, but it is designed for speech recognition and segmented into individual utterances, making it unsuitable for dialogue modeling.
3 Corpus Construction Methodology
Figure 1: Corpus construction methodology proposed in this work.
Figure 1 shows the overall flow of the corpus construction process. To construct a target-language (in our case, Japanese) spoken dialogue corpus, we collected data from the internet, excluded inappropriate content from the linguistic, dialogue, and acoustic perspectives by identifying the spoken language, extracting dialogue utterances, and removing background noise.
3.1 Data Collection
Since previous research Seki et al. (2023) has demonstrated that a diverse range of speech can be collected from YouTube, we used it as one of our data sources. We searched YouTube with randomly extracted keywords from Wikipedia page names, as in the previous research Takamichi et al. (2021), resulting in approximately 600k audio files totaling around 180k hours of audio data.
YouTube contains not only dialogue-dominant videos but also non-speech videos like music and monologue videos like game commentaries. This resulted in a low proportion of dialogue data, making it challenging to secure a sufficient amount of data. To address this, we also collected data from podcasts to expand the scale of our dataset. Podcasts are speech platforms, making it efficient to gather speech data from them. Furthermore, PodcastIndex [^2] https://podcastindex.org/ provides extensive metadata, including labels that indicates which language the content is. We retrieved RSS feed URLs of all podcast stations labeled as Japanese from podcast Index. Subsequently, we searched and downloaded for any audio URLs listed on the collected RSS feeds. As a result, we obtained approximately 880k audio files totaling around 140k hours of audio data.
3.2 Data Selection
3.2.1 Extracting Japanese Speech Data
For filtering out non-Japanese data, we used Whisper’s language identification model Radford et al. (2023), with the probability $p$ of being Japanese. Then, we extracted data only if satisfying $p$ exceed $0.8$. This process extracted $55.7\%$ of the data from YouTube and $84.7\%$ of the data from podcasts.
3.2.2 Extracting Dialogue Speech Data
From this Japanese speech dataset, we specifically extracted dialogue speech data. This requires detecting dialogue segments within the audio data. Related to this, speaker diarization (SD) is studied to identify who is speaking and when. SD analyzes the audio data to detect speech segments and links segments spoken by the same speaker, outputting pairs of speech segments and speaker IDs. In this study, we used publicly available pre-trained SD models (PyAnnote, Plaquet and Bredin, 2023) to obtain speech segments with speaker IDs.
Based on the speech segments (hereafter referred to as turns), we split the audio data into separate dialogues at gaps of 5 seconds or more. Next, we filtered out dialogues where a single speaker’s turns account for more than $80\%$ of the time, ensuring that only valid conversations are selected. This is to consider dialogues dominated by a single speaker as monologues. Here, we allowed dialogues with two or more distinct speaker IDs as valid types of conversations.
Through the above process, we obtained pairs of dialogue speech data and their labels (each turn’s duration and speaker ID). The proportion of Japanese speech data containing dialogues was $41.9\%$ for the YouTube data and approximately $45.0\%$ for the podcast data.
3.3 Data Cleansing
Podcasts often include background music (BGM), which acts as noise for speech generation models, so it needs to be removed. Techniques for extracting speech from data mixed with BGM are studied in the fields of speech enhancement and source separation, and recently, machine learning models have achieved high performance in this area. We applied a pre-trained speech enhancement model (Demucs Rouard et al., 2023) for data cleansing to all the collected audio and obtained J-CHAT.
4 Corpus Analysis
4.1 Dataset Size
Table 2: Corpus statistics by its subsets, YouTube and Podcast. # means “number of”.
feature YouTube Podcast Total
total duration[hr] 11,001 57,891 68,892
# dialogue 1,013,488 3,924,009 4,937,497
mean duration [s] 39.07 53.11 50.23
mean # turns 7.58 10.68 10.10
mean # speakers 3.23 3.12 3.14
Table 2 shows the statistical analysis of J-CHAT. J-CHAT consists of a total of 69k hours of Japanese speech data, with the YouTube subset accounting for 11k hours and the podcast subset for 58k hours.
A notable difference between the subsets is that the average duration of dialogues in the podcast subset is approximately 1.4 times longer, which is attributed to a higher number of turns per dialogue. On the other hand, the average number of speakers per dialogue is nearly the same for both subsets.
4.2 Phonetic Diversity
Figure 2: Distribution of HuBERT Hsu et al. (2021) features extracted from J-CHAT (ours), STUDIES (simulated dialogue), and JNV (non-verbal expression).
To confirm the phonetic diversity of J-CHAT, we analyzed the distribution of HuBERT Hsu et al. (2021) features. HuBERT is a speech encoding model similar to BERT Devlin et al. (2019) which captures phonetic information from speech Hsu et al. (2021). We randomly sampled 1000 frame-wise HuBERT features from each subset of J-CHAT, as well as from STUDIES Saito et al. (2022) and JNV Xin et al. (2024b). For each sampled point, we compute HuBERT features by randomly selecting 5-second intervals. These features were then reduced in dimensionality using t-SNE Van der Maaten and Hinton (2008), as illustrated in Figure 2. The distributions for STUDIES and JNV, which correspond to text-aligned speech and non-verbal expressions respectively, were found to be limited to specific regions. In contrast, the subsets of J-CHAT encompassed both regions, illustrating the broad phonetic diversity present in J-CHAT.
5 Experiments
To validate that J-CHAT corpus is a large-scale corpus suitable for training dialogue-oriented SLMs, we trained and evaluated the performance of dialogue generative spoken language model (dGSLM, Nguyen et al., 2023). dGSLM follows a framework proposed by Lakhotia et al., 2021 which consists of three distinct modules: speech-to-unit, unit language model, and unit-to-speech (vocoder).
Our experiments include three models and resynthesis samples for comparison: a resynthesized J-CHAT test subset utilizing speech-to-unit and vocoder (resynth), dGSLM on the YouTube subset of J-CHAT (dGSLM-YouTube), dGSLM trained on the podcast subset (dGSLM-podcast), and dGSLM trained on all subsets of J-CHAT (dGSLM-J-CHAT). Trained models, generated samples and training data are available1.
5.1 Experimental Conditions
We split the dialogues in the J-CHAT corpus into two channels by swapping the output channel of speech according to the turn-taking event. Subsequently we randomly split the each subset of J-CHAT corpus into train/valid/test sets. Then, we performed discretization of speech using k-means clustering on HuBERT-extracted features using train set from both subsets. The number of clusters for k-means was set to 1000.
For the vocoder used for speech generation from the discretized speech, we used HiFi-GAN Kong et al. (2020) conditioned with speaker information from XVector Snyder et al. (2018), as used in previous work Kharitonov et al. (2022). Please refer to appendix A for more details.
5.2 Evaluation
For evaluation, we performed subjective listening tests on the mean opinion score (MOS) on the naturalness and meaningfulness of dialogues following previous work Nguyen et al. (2023). For each subjective listening test, we recruited 60 Japanese native speakers, and each listener evaluated 8 samples. For sample generation used in the evaluation, we used the first 5 seconds of the test set derived from J-CHAT to prompt the model, then performed inference to predict the next 25 seconds of the dialogue based on those initial 5 seconds. For the sampling method, we used the beam search with a beam size of 5. When synthesizing speech from discritized speech, we conditioned the vocoder with JVS001 (male) and JVS002 (female) speaker from the JVS corpus Takamichi et al. (2020). This resulted in the male speaker’s voice being heard from the first audio channel and the female speaker’s voice from the second channel.
5.3 Result
Table 3: Result for MOS tests with their 95% confidence intervals.
Model Naturalness Meaningfulness
resynth 2.55 $\pm$ 0.18 2.48 $\pm$ 0.18
dGSLM-Youtube 1.44 $\pm$ 0.13 1.56 $\pm$ 0.14
dGSLM-podcast 1.44 $\pm$ 0.13 1.52 $\pm$ 0.13
dGSLM-J-CHAT 2.28 $\pm$ 0.19 2.18 $\pm$ 0.19
Table 3 shows the results of the subjective evaluation. From these results, it can be seen that dGSLM-J-CHAT achieves the best performance among the dGSLM generated samples in both naturalness and meaningfulness. This suggests that J-CHAT is a useful corpus for constructing generative dialogue language models. We can also see that there is no statistical significance between dGSLM-YouTube and dGSLM-podcast, despite dGSLM-podcast being trained on a dataset approximately four times larger in number of dialogues. This indicates that simply scaling up the dataset is not enough to enhance dGSLM performance. However, there is a significant difference between resynth and dGSLM-J-CHAT in terms of both naturalness and meaningfulness. The trained model occasionally produces sensible words, but the generated dialogue often lacks coherence. This might be improved with better modeling or by increasing the dataset size. For a detailed explanation on the statistical significance of each result, refer to the appendix B.2.
6 Conclusion
We have demonstrated a scalable dataset construction method. The experimental results showed that collecting from multiple domains, such as YouTube and podcasts, is beneficial in training dialogue language models. We hope that this corpus will contribute to the advancement of research in dialogue speech models.
Limitations
In this study, we believe that obtaining data from multiple domains, such as YouTube and podcasts, allows us to gather a wide range of speech. However, we also recognize that there are limitations to the range that can be covered. For example, OGVC corpus Arimoto et al. (2012) reports a method to collect a laughter corpus from the internet, but found that the amount of screaming was limited because the internet data is typically collected under controlled conditions. Similarly, podcast dialogues are intended for listeners, so informal expressions used among close friends are likely to be limited.
Ethics Statement
We recognize that speech synthesis technology carries the risk of being misused for unauthorized generation of others’ voices, commonly referred to as voice cloning. To prevent such misuse, we converted the collected speech into latent variables known as HuBERT features, which lack speaker identity information. Although HuBERT does not completely eliminate speaker information, it obscures it sufficiently to make the restoration of speaker identity challenging, thus serving as an effective countermeasure against voice cloning. In our experiments, we synthesized speech using the speaker identity of individuals who had consented to their voice being used for synthesis at the time of recording, ensuring that no unauthorized voice replication was performed.
On the subjective evaluation conducted on Section 5.2, participants are paid with adequate amount of money and they were instructed with the purpose of the research and how the collected data will be used. Please refer to the Section B for the details on the subjective evaluation.
Acknowledgements
This work was supported by AIST KAKUSEI project (FY2023), JST Moonshot JPMJMS2011 and JST FOREST JPMJFR226V.
Appendix A dGSLM Training Details
For the implementation of dGSLM, we used fairseq [^3] https://github.com/facebookresearch/fairseq repository by Meta which includes official implementation of the model. This repository is licensed as MIT.
The number of train/valid/test is 1,003,397/9,991/100 dialogues for YouTube subset and for podcast subset it is 3,885,007/38,902/100 dialogues.
For HuBERT model used in the speech-to-unit, we used the pretrained Japanese HuBERT model Sawada et al. (2024) [^4] https://huggingface.co/rinna/japanese-hubert-base. For k-means clustering, we used the original implementation of the dGSLM except for the number of clusters which is set to 1,000.
The dGSLM model was configured as in the previous research Nguyen et al. (2023). The number of model parameters in each dGSLM model is 23,638,529.
For training dGSLM model, we used 32 NVIDIA V100 GPUs with the learning rate of $2\times 10^{-4}$. The training was performed for 100,000 steps. The total duration of training was around 48 hours. The batch size was 36,864 tokens and the the maximum number of tokens per sample are set to 3,000. This is equivalent to the 1 minute of dialogue. For other hyper-parameters regarding the training, we followed the original paper Nguyen et al. (2023).
For the XVector conditioned vocoder, we used the pretrained XVector model [^5] https://huggingface.co/speechbrain/spkrec-xvect-voxceleb. The vocoder was trained with the JVS Takamichi et al. (2020) and JVNV Xin et al. (2024a) corpus. These two corpora include Japanese reading-style, studio-quality speech without/with non-verbal expression, respectively. Training of vocoder took 2 days with 4 NVIDIA V100 GPUs.
Appendix B Subjective Evaluation of the dGSLM Outputs
B.1 Participant Details
Participants of the each subjective evaluation were recruited on https://lancers.jp. They are paid with 75 Japanese yen upon the completion of the test. Given that the minimum wage of Japan is 1,004 Japanese yen and the test took approximately 4 minutes, the payment is adequate.
Before the participation, the participants were instructed with the research purpose and how the collected data will be used in this research. The data collection protocol is approved by our institution’s ethics review board.
Figure 3: Screenshot of instruction given to the participants on the naturalness MOS evaluation.
Figure 4: Screenshot of instruction given to the participants on the meaningfulness MOS evaluation.
Figures  3 and 4 show the screenshots of the instructions given to the participants for the naturalness and meaningfulness MOS evaluations, respectively. The English translation of the instruction for the naturalness MOS evaluation is as follows:
Listen to the audio and rate the naturalness of the sound on a scale of 1 to 5 (very unnatural (1) to very natural (5)). Please conduct the evaluation using headphones in a quiet environment. Please note that some audio samples may have degraded sound quality, so please be careful with the volume level. Do not use your browser’s “back” or “reload” buttons during the test.
The English translation of the instruction for the meaningfulness MOS evaluation is as follows:
Listen to the audio and rate the meaningfulness of the dialogue on a 5-point scale from very meaningless (1) to very meaningful (5). Please evaluate in a quiet environment using headphones. Please note that some of the audio samples may have degraded sound quality, so please be careful with the volume. Please do not use your browser’s “back” or “reload” buttons during the test.
B.2 P-values from the Subjective Evaluation
Figure 5: P-values from the subjective evaluation on the naturalness of the dGSLM output samples
Figure 6: P-values from the subjective evaluation on the meaningfulness of the dGSLM output samples
Figure 5 and Figure 6 shows the p-values from the subjective evaluation on the dGSLM output conducted on Section 5.2. Specifically, we performed two sided T-test on the MOS score.

### v2

@@
 Abstract
-Spoken dialogue plays a crucial role in human-AI interactions, necessitating dialogue-oriented spoken language models (SLMs).
-To develop versatile SLMs, large-scale and diverse speech datasets are essential.
-Additionally, to ensure hiqh-quality speech generation, the data must be spontaneous like in-wild data and must be acoustically clean with noise removed.
-Despite the critical need, no open-source corpus meeting all these criteria has been available.
-This study addresses this gap by constructing and releasing a large-scale spoken dialogue corpus, named Japanese Corpus for Human-AI Talks (J-CHAT), which is publicly accessible.
-Furthermore, this paper presents a language-independent method for corpus construction and describes experiments on dialogue generation using SLMs trained on J-CHAT.
-Experimental results indicate that the collected data from multiple domains by our method improve the naturalness and meaningfulness of dialogue generation.
-J-CHAT: Japanese Large-scale Spoken Dialogue Corpus
-for Spoken Dialogue Language Modeling
-Wataru Nakata1∗  Kentaro Seki1 thanks:  Equal contribution.
-Hitomi Yanaka1 Yuki Saito1Shinnosuke Takamichi1,2Hiroshi Saruwatari1 1 The University of Tokyo, Japan  2 Keio University, Japan nakata-wataru855@g.ecc.u-tokyo.ac.jp
-1 Introduction
-Spoken dialogue systems have garnered significant attention for their role in enabling human-AI communication Mitsui et al.
+Spoken dialogue is essential for human-AI interactions, providing expressive capabilities beyond text.
+Developing effective spoken dialogue systems (SDSs) requires large-scale, high-quality, and diverse spoken dialogue corpora.
+However, existing datasets are often limited in size, spontaneity, or linguistic coherence.
+To address these limitations, we introduce J-CHAT, a 76,000-hour open-source Japanese spoken dialogue corpus.
+Constructed using an automated, language-independent methodology, J-CHAT ensures acoustic cleanliness, diversity, and natural spontaneity.
+The corpus is built from YouTube and podcast data, with extensive filtering and denoising to enhance quality.
+Experimental results with generative spoken dialogue language models trained on J-CHAT demonstrate its effectiveness for SDS development.
+By providing a robust foundation for training advanced dialogue models, we anticipate that J-CHAT will drive progress in human-AI dialogue research and applications.
+Keywords: spoken dialogue corpus, speech datasets, Japanese, spoken language modeling, language resources
+[…]
+J-CHAT: Japanese Large-scale Spoken Dialogue Corpus for Spoken Dialogue Language Modeling
+Wataru Nakata1∗, Kentaro Seki1∗, Hitomi Yanaka1
+Yuki Saito1 Shinnosuke Takamichi1,2, Hiroshi Saruwatari1
+1The University of Tokyo, Japan
+2Keio University, Japan
+nakata-wataru855@g.ecc.u-tokyo.ac.jp
+Equal contribution.
+Abstract content
+1.  Introduction
+To realize human-AI interaction through spoken language, the development of spoken dialogue systems (SDSs) is crucial.
+Traditional SDSs have been constructed by cascading multiple modules, such as speech recognition, response generation, and speech synthesis Huang et al.
 (2024).
-Spoken language holds rich expressiveness through tone, intonation, and non-verbal vocalizations such as laughter and screams, highlighting the importance of spoken dialogue systems.
-To leverage the expressive capabilities of spoken language in spoken dialogue systems, it is essential to develop spoken language models (SLMs) that model the linguistic structures of spoken dialogue.
-In other words, SLMs are crucial for facilitating smooth and effective human-AI communication.
-Generative spoken language modeling (GSLM, Lakhotia et al., 2021) using HuBERT Hsu et al.
-(2021) as a speech feature extractor is a well-known example of SLMs.
-Training SLMs requires a large-scale speech corpus.
-Open-source corpora are crucial because they lower research barriers and ensure reproducible results.
-Despite their importance, the development of spoken dialogue corpora has been limited.
-Current methods for constructing spoken dialogue corpora Cieri et al.
-(2004); Saito et al.
-(2022) are labor-intensive making it challenging to scale the corpora.
-Furthermore, large-scale spoken dialogue corpora have primarily been constructed for English Cieri et al.
-(2004), with limited progress in other languages, such as Japanese.
-Therefore, we construct and release a new open-source corpus named “Japanese Corpus for Human-AI Talks” (J-CHAT) [^1] https://huggingface.co/datasets/sarulab-speech/J-CHAT.
-To collect large-scale speech data, we propose a fully automated thus scalable method for corpus construction, including data collection from the web.
-Gathering in-the-wild data ensures the collection of spontaneous speech covering a wide range of conversational topics.
-Additionally, we use noise removal techniques to make the data acoustically clean.
-Our contributions include:
-  • Constructing and releasing a novel, open-source, and large-scale speech corpus.
-  • Providing a spontaneous and clean speech corpus, thereby facilitating the generation of natural dialogues.
-  • Developing a corpus construction method that is independent of the language.
+However, this cascaded SDS approach struggles to account for subtle nuances and nonverbal vocal expressions, which are often lost in transcription yet are critical elements that distinguish spoken dialogue from text-based chat.
+Recently, it has been demonstrated that end-to-end SDSs can be realized using machine learning methods trained on large-scale data, and several such methods have been proposed Nguyen et al.
+(2023); Mitsui et al.
+(2023).
+This approach is gaining attention because it is expected to enable smoother human-AI interaction by fully leveraging the information conveyed through spoken language.
+The realization of end-to-end SDS requires large-scale datasets.
+For instance, previous studies on synthesizing dialogue speech using end-to-end SDS have utilized 20k hours of speech Nguyen et al.
+(2023) and 100k hours of speech Borsos et al.
+(2023), suggesting that tens of thousands of hours of dialogue speech corpora are necessary to achieve this approach.
+In addition to size, the quality and diversity of the corpus are also crucial.
+In dialogue systems, since the final output is speech, it is desirable for the training data to be clean and free of noise.
+Furthermore, machine learning models tend to degrade in performance when applied to out-of-domain data.
+Therefore, it is desirable to have a diverse dataset, particularly including spontaneous speech, which is often not covered by existing studio-recorded corpora.
+However, large-scale open-source dialogue speech corpora remain scarce, and methods for their construction are underdeveloped.
+The largest available resource, Seamless Interaction Agrawal et al.
+(2025), contains over 4,000 hours of dialogue speech.
+While this corpus is sufficiently large for training SDSs, its manual recording under controlled conditions makes it financially impractical to reproduce in other languages.
+Although several studies have investigated automatic methods for constructing large-scale speech corpora in recent years, these efforts have primarily targeted speech recognition Chen et al.
+(2021); Takamichi et al.
+(2021) or speech synthesis Seki et al.
+(2023), typically segmenting speech into short utterance-level units.
+Consequently, such corpora lack key characteristics of dialogue, including linguistic coherence across turns and natural turn-taking dynamics.
+To date, no general method for automatically constructing a large-scale dialogue speech corpus has been explored.
+To address this issue, we propose a method for constructing large-scale dialogue speech corpora and have publicly released the constructed corpus, named “Japanese Corpus for Human-AI Talks” (J-CHAT) [^1] https://huggingface.co/datasets/sarulab-speech/J-CHAT.
+The corpus is distributed under CC BY-NC 4.0 license for “information analysis” which is defined in Japanese copyright act article 30-4222https://laws.e-gov.go.jp/law/345AC0000000048#Mp-Ch_2-Se_3-Ss_5-At_30_4.
+Our corpus construction method automatically filters and collects data on a dialogue-unit basis from the internet.
+Since the entire process is automated, it is highly scalable.
+The constructed J-CHAT corpus contains 76k hours of Japanese dialogue speech data, which is comparable to the existing large-scale corpora for the development of end-to-end SDSs.
+Additionally, background music removal procedures ensure the acoustic cleanliness of the dataset.
+Moreover, by collecting wild data from multiple domains, the corpus includes spontaneous speech and consists of diverse data.
+Our contributions are as follows:
+  • We propose a method for constructing a large-scale corpus for end-to-end SDSs. Our proposed method constructs a corpus that is acoustically clean, diverse, and includes spontaneous speech.
+  • Our proposed method is an automated and language-independent approach, making it easy to construct corpora for other languages.
+  • We constructed and released the Japanese dialogue speech corpus, J-CHAT. Also, we experimentaly validated its effectiveness for end-to-end SDSs.
 Table 1: Comparison of speech corpora related to this study, sorted by size.
 Corpus name Size(hours) Open-source Dialogue Spontaneous Clean speech
-STUDIES (Saito et al., 2022) $8.2$
-& ✅   ✅ DailyTalk (Lee et al., 2023) $20$ Fisher (Cieri et al., 2004) $2k$   ✅ ✅ ✅ GigaSpeech (Chen et al., 2021) $33k$   -   J-CHAT (This study) $69k$
-2 Related Work
-2.1 Spoken Language Models
-Large-scale SLMs have been reported to enhance the performance of various downstream tasks.
-Training these models requires corpora that are significantly larger than traditional ones.
-For instance, research using 100k hours of audio for dialogue generation Borsos et al.
-(2023), 680k hours for speech recognition Radford et al.
-(2023), and 4.5M hours for speech translation Barrault et al.
-(2023) has been reported.
-These studies demonstrate that increasing the amount of data leads to performance improvements, motivating us to construct large-scale speech corpora.
-However, these studies primarily focus on model architecture improvements, and the datasets used in the experiments are often not publicly available, with little to no detail on their construction methods.
-One research problem that suffers from a lack of data is dialogue spoken language modeling Nguyen et al.
-(2023).
-In a previous study, 2000 hours of dialogue data were used, drawing from one of the largest available dialogue speech data resources.
-Despite this substantial dataset, the researchers reported that its size was insufficient for effective dialogue spoken language modeling.
-By proposing and releasing a method for constructing a large-scale corpus specifically for the dialogue domain, we aim to facilitate discussions and advancements in corpus construction techniques.
-2.2 Existing Speech Corpora
-Training SLMs for dialogues requires large-scale, spontaneous and acoustically clean dialogue speech data.
-Table 1 shows corpora related to this study.
-STUDIES corpus Saito et al.
-(2022) and DailyTalk corpus Lee et al.
-(2023) are open-source dialogue corpora designed for multi-turn conversations, but they are very small in size because they are manually recorded spoken dialogues.
-Fisher corpus Cieri et al.
-(2004) is more extensive but is not open-source.
-GigaSpeech corpus Chen et al.
-(2021) is a large-scale open-source corpus, but it is designed for speech recognition and segmented into individual utterances, making it unsuitable for dialogue modeling.
-3 Corpus Construction Methodology
+STUDIES Saito et al. (2022) $8.2$
+$\checkmark$   $\checkmark$ DailyTalk Lee et al.
+(2023) 2020 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ CallHome-JP Alexandra and Zipperlen (1996) $49$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ MultiDialog Park et al.
+(2024) $340$ $\checkmark$ $\checkmark$ $\checkmark$   LibriTTS Zen et al.
+(2019) $585$ $\checkmark$     $\checkmark$ SSSD Sheikh et al.
+(2025) $727$ $\checkmark$ $\checkmark$   $\checkmark$ Fisher Cieri et al.
+(2004) $2k$   $\checkmark$ $\checkmark$ $\checkmark$ Seamless Interaction Agrawal et al.
+(2025) $4k$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ GigaSpeech Chen et al.
+(2021) $33k$ $\checkmark$   $\checkmark$   J-CHAT (This study) $76k$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 Figure 1: Corpus construction methodology proposed in this work.
-Figure 1 shows the overall flow of the corpus construction process.
-To construct a target-language (in our case, Japanese) spoken dialogue corpus, we collected data from the internet, excluded inappropriate content from the linguistic, dialogue, and acoustic perspectives by identifying the spoken language, extracting dialogue utterances, and removing background noise.
-3.1 Data Collection
-Since previous research Seki et al.
-(2023) has demonstrated that a diverse range of speech can be collected from YouTube, we used it as one of our data sources.
-We searched YouTube with randomly extracted keywords from Wikipedia page names, as in the previous research Takamichi et al.
-(2021), resulting in approximately 600k audio files totaling around 180k hours of audio data.
+2.  Related Work
+A comparison between existing corpora and our J-CHAT corpus is shown in Table 1.
+The STUDIES corpus Saito et al.
+(2022) is an open-source conversational speech corpus; however, since it consists of scripted speech recorded in a studio, it does not include spontaneous speech and covers only limited domains.
+The DailyTalk corpus (Lee et al., 2023) is also an open-source conversational speech corpus, but it contains only about 20 hours of data, which is relatively small in scale.
+The CallHome Japanese subset (hereinafter referred to as “CallHome-JP”) is the largest spontaneous conversational speech dataset in Japanese.
+Although it is open-source, its scale remains insufficient for training end-to-end SDSs.
+The MultiDialog corpus Park et al.
+(2024), and the SSSD Sheikh et al.
+(2025) corpus are English spontaneous conversational speech datasets, but their size are still insufficient for end-to-end SDS.
+The Fisher corpus Cieri et al.
+(2004) is a sufficiently large conversational speech dataset, but it is not open-source, making experimental reproducibility challenging.
+The LibriTTS corpus Zen et al.
+(2019) and GigaSpeech corpus Chen et al.
+(2021) are open-source corpora; however, as it is designed for speech synthesis and recognition, respectively, its data is segmented at the utterance level, lacking conversational context-specific information.
+Recently, Seamless Interaction Agrawal et al.
+(2025) has been released which is over 4k hours in duration making it a viable training dataset for SDSs.
+However, their corpus construction methodology was based on manual recording in a controlled environment limiting the scalability required to reproduce SDSs in other languages.
+In contrast, our J-CHAT corpus was collected from in-the-wild sources, allowing for the inclusion of diverse speech.
+Furthermore, using in-the-wild sources allows scalability of the dataset size making the construction SDSs in many languages possible.
+With a scale of 76k hours, which is over 1,000 times larger than CallHome-JP, it provides a dataset large enough for the development of end-to-end SDSs.
+Furthermore, as an open-source resource, it is expected to facilitate research and development in SDSs.
+Unlike utterance-level segmentation, J-CHAT is segmented at the dialogue level, which enables modeling of discourse phenomena such as contextual coherence and turn-taking behavior.
+3.  Corpus Construction Methodology
+Figure 1 illustrates the overall workflow of our corpus construction process.
+To build a spoken dialogue corpus for a target language (Japanese in this study), we collected audio data from the internet.
+We then removed inappropriate content by performing language identification, dialogue extraction, and background-noise removal.
+Finally, we transcribed the spoken content using an automatic speech recognition (ASR) model.
+3.1.  Data Collection
+Since previous research Takamichi et al.
+(2021) has demonstrated that a diverse range of speech can be collected from YouTube, we used it as one of our primary data sources.
+We searched YouTube using randomly selected Wikipedia page titles as keywords, following previous research (Takamichi et al., 2021), which resulted in approximately 600k audio files totaling about 180k hours of speech.
 YouTube contains not only dialogue-dominant videos but also non-speech videos like music and monologue videos like game commentaries.
@@
 Podcasts are speech platforms, making it efficient to gather speech data from them.
-Furthermore, PodcastIndex [^2] https://podcastindex.org/ provides extensive metadata, including labels that indicates which language the content is.
+Furthermore, PodcastIndex [^3] https://podcastindex.org/ provides extensive metadata, including labels that indicate which language the content is.
 We retrieved RSS feed URLs of all podcast stations labeled as Japanese from PodcastIndex.
@@
 As a result, we obtained approximately 880k audio files totaling around 140k hours of audio data.
-3.2 Data Selection
-3.2.1 Extracting Japanese Speech Data
-For filtering out non-Japanese data, we used Whisper’s language identification model Radford et al.
-(2023), with the probability $p$ of being Japanese.
-Then, we extracted data only if satisfying $p$ exceed $0.8$.
-This process extracted $55.7\%$ of the data from YouTube and $84.7\%$ of the data from podcasts.
-3.2.2 Extracting Dialogue Speech Data
-From this Japanese speech dataset, we specifically extracted dialogue speech data.
-This requires detecting dialogue segments within the audio data.
-Related to this, speaker diarization (SD) is studied to identify who is speaking and when.
+3.2.  Data Selection
+3.2.1.  Extracting Japanese Speech Data
+To filter out non-Japanese audio, we used Whisper’s language-identification model (Radford et al., 2023) and retained segments only when the probability of Japanese speech (p) exceeded 0.8.
+This process retained $55.7\%$ of the YouTube data and $84.7\%$ of the podcast data.
+3.2.2.  Extracting Dialogue Speech Data
+From this Japanese speech dataset, we specifically extracted dialogue segments.
+This requires detecting conversation regions within the audio.
+Speaker diarization (SD) techniques are commonly used to identify who is speaking and when.
 SD analyzes the audio data to detect speech segments and links segments spoken by the same speaker, outputting pairs of speech segments and speaker IDs.
-In this study, we used publicly available pre-trained SD models (PyAnnote, Plaquet and Bredin, 2023) to obtain speech segments with speaker IDs.
+In this study, we used a publicly available pre-trained SD model, PyAnnote Plaquet and Bredin (2023), to obtain speech segments with speaker IDs.
 Based on the speech segments (hereafter referred to as turns), we split the audio data into separate dialogues at gaps of 5 seconds or more.
 Next, we filtered out dialogues where a single speaker’s turns account for more than $80\%$ of the time, ensuring that only valid conversations are selected.
-This is to consider dialogues dominated by a single speaker as monologues.
+This treats dialogues dominated by a single speaker as monologues.
 Here, we allowed dialogues with two or more distinct speaker IDs as valid types of conversations.
-Through the above process, we obtained pairs of dialogue speech data and their labels (each turn’s duration and speaker ID).
+Through the above process, we obtained pairs of dialogue-speech data and their labels (each turn’s duration and speaker ID).
 The proportion of Japanese speech data containing dialogues was $41.9\%$ for the YouTube data and approximately $45.0\%$ for the podcast data.
-3.3 Data Cleansing
-Podcasts often include background music (BGM), which acts as noise for speech generation models, so it needs to be removed.
-Techniques for extracting speech from data mixed with BGM are studied in the fields of speech enhancement and source separation, and recently, machine learning models have achieved high performance in this area.
-We applied a pre-trained speech enhancement model (Demucs Rouard et al., 2023) for data cleansing to all the collected audio and obtained J-CHAT.
-4 Corpus Analysis
-4.1 Dataset Size
+$3.$  Data Cleansing
+YouTube and podcasts often include background music (BGM), which acts as noise for speech generation models, so it needs to be removed.
+Techniques for extracting speech from data mixed with BGM are studied in the fields of speech enhancement and source separation, and recently, machine learning models have achieved high performance in this area Koizumi et al.
+(2023); Rouard et al.
+(2023).
+We applied a pre-trained speech enhancement model Demucs Rouard et al.
+(2023) for data cleansing to all the collected audio and obtained the J-CHAT corpus.
+3.4.  ASR
+Some SDSs require transcription of the speech for training Défossez et al.
+(2024).
+To meet this need, we created the transcription using ASR model.
+For ASR model, we used reazonspeech-nemo-v2444https://huggingface.co/reazon-research/reazonspeech-nemo-v2.
+For each subword in transcript, the alignment information is also provided.
+4.  Corpus Analysis
+4.1.  Dataset Size
 Table 2: Corpus statistics by its subsets, YouTube and Podcast. # means “number of”.
 feature YouTube Podcast Total
-total duration[hr] 11,001 57,891 68,892
-# dialogue 1,013,488 3,924,009 4,937,497
+total duration[hr] 11,017 65,019 76,036
+# dialogue 1,015,109 4,409,405 5,424,514
 mean duration [s] 39.07 53.11 50.23
@@
 mean # speakers 3.23 3.12 3.14
-Table 2 shows the statistical analysis of J-CHAT.
-J-CHAT consists of a total of 69k hours of Japanese speech data, with the YouTube subset accounting for 11k hours and the podcast subset for 58k hours.
+Table 2 shows the statistical analysis of J-CHAT corpus.
+J-CHAT consists of a total of 76k hours of Japanese speech data, with the YouTube subset accounting for 11k hours and the podcast subset for 65k hours.
+The corpus also provides predefined train/valid/test/other splits for each YouTube and Podcast subsets for the reproducibility of research.
+For the YouTube subset, the durations of the train/valid/test/other splits are 10872.5/108.7/1.2 hours respectively.
+For the Podcast subset the duration of train/valid/test/other splits are 57291.5/575/1.3 hours respectively.
 A notable difference between the subsets is that the average duration of dialogues in the podcast subset is approximately 1.4 times longer, which is attributed to a higher number of turns per dialogue.
 On the other hand, the average number of speakers per dialogue is nearly the same for both subsets.
-4.2 Phonetic Diversity
-Figure 2: Distribution of HuBERT Hsu et al. (2021) features extracted from J-CHAT (ours), STUDIES (simulated dialogue), and JNV (non-verbal expression).
-To confirm the phonetic diversity of J-CHAT, we analyzed the distribution of HuBERT Hsu et al.
-(2021) features.
-HuBERT is a speech encoding model similar to BERT Devlin et al.
-(2019) which captures phonetic information from speech Hsu et al.
-(2021).
-We randomly sampled 1000 frame-wise HuBERT features from each subset of J-CHAT, as well as from STUDIES Saito et al.
-(2022) and JNV Xin et al.
-(2024b).
-For each sampled point, we compute HuBERT features by randomly selecting 5-second intervals.
-These features were then reduced in dimensionality using t-SNE Van der Maaten and Hinton (2008), as illustrated in Figure 2.
-The distributions for STUDIES and JNV, which correspond to text-aligned speech and non-verbal expressions respectively, were found to be limited to specific regions.
-In contrast, the subsets of J-CHAT encompassed both regions, illustrating the broad phonetic diversity present in J-CHAT.
-5 Experiments
-To validate that J-CHAT corpus is a large-scale corpus suitable for training dialogue-oriented SLMs, we trained and evaluated the performance of dialogue generative spoken language model (dGSLM, Nguyen et al., 2023).
-dGSLM follows a framework proposed by Lakhotia et al., 2021 which consists of three distinct modules: speech-to-unit, unit language model, and unit-to-speech (vocoder).
+4.2.  Acoustic Cleanliness
+To ensure that the J-CHAT corpus is not significantly affected by noise, we evaluated its acoustic cleanliness using NISQA Mittag et al.
+(2021) and compared it with CallHome-JP (telephone recordings) and STUDIES (studio recordings).
+The average quality scores were as follows: $4.01$ for STUDIES, $1.98$ for CallHome-JP, $2.37$ for the YouTube subset of J-CHAT, and $2.99$ for the Podcast subset of J-CHAT.
+Although there is a quality gap between J-CHAT and the studio-recorded STUDIES corpus, J-CHAT’s quality is superior to that of CallHome-JP, the largest spontaneous dialogue corpus for Japanese.
+Furthermore, the Podcast subset of J-CHAT achieved a higher score than the YouTube subset, indicating the effectiveness of using podcasts as a data source for constructing dialogue corpora.
+This suggests that the noise level in the Podcast subset is sufficiently low for use in speech synthesis.
+4.3.  Dataset Diversity
+Figure 2: t-SNE visualization of sentence embeddings extracted from the ASRed transcripts from Podcast and YouTube subset of J-CHAT, STUDIES Saito et al.
+(2022) and CallHome-JP Alexandra and Zipperlen (1996) corpus.
+For each corpus, random 120 dialouges are extracted for calculation.
+Linguistic Diversity: As a dialogue corpus, J-CHAT is expected to cover a wide range of topics.
+To analyze its topic diversity, we compared the distribution of sentence embeddings from J-CHAT transcriptions with those from STUDIES (scripted) and CallHome-JP (spontaneous).
+We randomly selected $120$ dialogues from the YouTube subset of J-CHAT, and the Podcast subset of J-CHAT as well as the all dialogues in STUDIES.
+For CallHome-JP, we extracted all $120$ dialogues.
+Subsequently sentence embeddings are extracted using a pre-trained sentence embedding model [^5] https://huggingface.co/sonoisa/sentence-bert-base-ja-mean-tokens-v2.
+The t-SNE plot of the embeddings is shown in Figure 2.
+The results indicate that each subset of J-CHAT exhibits a broader distribution compared to CallHome-JP.
+Note that, although the figure based on this comparison at the same scale may make J-CHAT appear sparse, in reality, J-CHAT is significantly larger than CallHome-JP, ensuring that the region covered by CallHome-JP also contains a sufficient amount of data.
+Additionally, the distribution of STUDIES is distinct from those of CallHome-JP and J-CHAT because the dialogue topics of STUDIES are limited in conversations between a teacher and students in school.
+Furthermore, to conduct a quantitative analysis, we calculated the average pairwise cosine similarity for each dataset.
+The values for CallHome-JP, STUDIES, the YouTube subset, and the Podcast subset are $0.6164$, $0.5186$, $0.2390$, and $0.3457$, respectively.
+Since a lower cosine similarity indicates lower similarity, these results quantitatively demonstrate that the J-CHAT corpus covers a more diverse range of topics.
+Figure 3: Distribution of HuBERT Hsu et al. (2021) features extracted from J-CHAT (ours), STUDIES (simulated dialogue), and CallHome-JP (real dialogue).
+Phonetic Diversity: To demonstrate that J-CHAT includes spontaneous speech, we conducted an analysis of acoustic features to confirm that J-CHAT contains expression characteristics of spontaneous speech.
+In this analysis, we used HuBERT Hsu et al.
+(2021) features, which are expected to capture phonetic information, and compared them with the STUDIES corpus, which consists of scripted dialogue speech, and CallHome-JP, which consists of spontaneous dialogue speech.
+We randomly sampled 1,000 frame-wise HuBERT features from each subset of J-CHAT, as well as from STUDIES Saito et al.
+(2022) and CallHome-JP Alexandra and Zipperlen (1996).
+For each sampled point, we computed HuBERT features by randomly selecting 5-second intervals.
+The t-SNE plot is shown in Figure 3.
+The distributions for STUDIES were found to be limited to specific regions.
+In contrast, the subsets of J-CHAT and CallHome-JP encompassed both regions.
+These results indicate that J-CHAT covers phonetic diversity in spontaneous dialogue speech.
+5.  Experiments
+To validate that J-CHAT corpus is a large-scale corpus suitable for training dialogue-oriented spoken language models, we trained and evaluated the performance of dialogue-generative spoken language model (dGSLM, Nguyen et al.
+(2023)).
+dGSLM follows a framework proposed by Lakhotia et al.
+(2021) which consists of three distinct modules: speech-to-unit, unit language model, and unit-to-speech (vocoder).
 Our experiments include three models and resynthesis samples for comparison: a resynthesized J-CHAT test subset utilizing speech-to-unit and vocoder (resynth), dGSLM on the YouTube subset of J-CHAT (dGSLM-YouTube), dGSLM trained on the podcast subset (dGSLM-podcast), and dGSLM trained on all subsets of J-CHAT (dGSLM-J-CHAT).
 Trained models, generated samples and training data are available1.
-5.1 Experimental Conditions
+5.1.  Experimental Conditions
 We split the dialogues in the J-CHAT corpus into two channels by swapping the output channel of speech according to the turn-taking event.
-Subsequently we randomly split the each subset of J-CHAT corpus into train/valid/test sets.
-Then, we performed discretization of speech using k-means clustering on HuBERT-extracted features using train set from both subsets.
-The number of clusters for k-means was set to 1000.
+We used the train/valid/test sets explained in Section 4.1.
+Then, we performed discretization of speech using k-means clustering on HuBERT-extracted features using train set from both YouTube and Podcast subsets.
+The number of clusters for k-means was set to 1,000.
 For the vocoder used for speech generation from the discretized speech, we used HiFi-GAN Kong et al.
@@
 (2022).
-Please refer to appendix A for more details.
-5.2 Evaluation
+5.2.  Training Details
+For the implementation of dGSLM, we used official implementation of the model [^6] https://github.com/facebookresearch/fairseq.
+For HuBERT model used in the speech-to-unit, we used the pretrained Japanese HuBERT model Sawada et al.
+(2024) [^7] https://huggingface.co/rinna/japanese-hubert-base.
+For k-means clustering, we used the original implementation of the dGSLM except for the number of clusters which was set to $1,000$.
+The dGSLM model was configured as in the previous research Nguyen et al.
+(2023).
+For training dGSLM model, we used 32 NVIDIA V100 GPUs with the learning rate of $2\times 10^{-4}$.
+The training was performed for $100,000$ steps.
+The batch size was $36,864$ tokens and the the maximum number of tokens per sample are set to $3,000$.
+This was equivalent to the $1$ minute of dialogue.
+For other hyper-parameters regarding the training, we followed the original paper Nguyen et al.
+(2023).
+Table 3: MOS test results with their 95% confidence intervals.
+Model Naturalness Meaningfulness
+resynth 2.55 $\pm$ 0.18 2.48 $\pm$ 0.18
+dGSLM-Youtube 1.44 $\pm$ 0.13 1.56 $\pm$ 0.14
+dGSLM-podcast 1.44 $\pm$ 0.13 1.52 $\pm$ 0.13
+dGSLM-J-CHAT 2.28 $\pm$ 0.19 2.18 $\pm$ 0.19
+For the XVector conditioned vocoder, we used the pretrained XVector model [^8] https://huggingface.co/speechbrain/spkrec-xvect-voxceleb.
+The vocoder was trained with the JVS Takamichi et al.
+(2020) and JVNV Xin et al.
+(2024) corpus.
+These two corpora include Japanese reading-style, studio-quality speech without/with non-verbal expression, respectively.
+Training of vocoder took 2 days with 4 NVIDIA V100 GPUs.
+5.3.  Evaluation
 For evaluation, we performed subjective listening tests on the mean opinion score (MOS) on the naturalness and meaningfulness of dialogues following previous work Nguyen et al.
@@
 (2020).
-This resulted in the male speaker’s voice being heard from the first audio channel and the female speaker’s voice from the second channel.
-5.3 Result
-Table 3: Result for MOS tests with their 95% confidence intervals.
-Model Naturalness Meaningfulness
-resynth 2.55 $\pm$ 0.18 2.48 $\pm$ 0.18
-dGSLM-Youtube 1.44 $\pm$ 0.13 1.56 $\pm$ 0.14
-dGSLM-podcast 1.44 $\pm$ 0.13 1.52 $\pm$ 0.13
-dGSLM-J-CHAT 2.28 $\pm$ 0.19 2.18 $\pm$ 0.19
+This resulted in the male speaker’s voice being heard from the first audio channel (left) and the female speaker’s voice from the second channel (right).
+5.4.  Result
 Table 3 shows the results of the subjective evaluation.
@@
 This might be improved with better modeling or by increasing the dataset size.
-For a detailed explanation on the statistical significance of each result, refer to the appendix B.2.
-6 Conclusion
-We have demonstrated a scalable dataset construction method.
-The experimental results showed that collecting from multiple domains, such as YouTube and podcasts, is beneficial in training dialogue language models.
-We hope that this corpus will contribute to the advancement of research in dialogue speech models.
-Limitations
-In this study, we believe that obtaining data from multiple domains, such as YouTube and podcasts, allows us to gather a wide range of speech.
-However, we also recognize that there are limitations to the range that can be covered.
-For example, OGVC corpus Arimoto et al.
-(2012) reports a method to collect a laughter corpus from the internet, but found that the amount of screaming was limited because the internet data is typically collected under controlled conditions.
-Similarly, podcast dialogues are intended for listeners, so informal expressions used among close friends are likely to be limited.
-Ethics Statement
-We recognize that speech synthesis technology carries the risk of being misused for unauthorized generation of others’ voices, commonly referred to as voice cloning.
-To prevent such misuse, we converted the collected speech into latent variables known as HuBERT features, which lack speaker identity information.
-Although HuBERT does not completely eliminate speaker information, it obscures it sufficiently to make the restoration of speaker identity challenging, thus serving as an effective countermeasure against voice cloning.
-In our experiments, we synthesized speech using the speaker identity of individuals who had consented to their voice being used for synthesis at the time of recording, ensuring that no unauthorized voice replication was performed.
-On the subjective evaluation conducted on Section 5.2, participants are paid with adequate amount of money and they were instructed with the purpose of the research and how the collected data will be used.
-Please refer to the Section B for the details on the subjective evaluation.
-Acknowledgements
+6.  Conclusion
+In this study, we presented J-CHAT, a large-scale Japanese spoken dialogue corpus, and proposed an automated and language-independent methodology for its construction.
+The experimental results demonstrated that training on diverse dialogue data from multiple data sources, such as YouTube and podcasts, significantly improves spoken dialogue systems models than collecting from single data source.
+Future work includes further improving dialogue modeling techniques to enhance the quality of generated speech.
+7.  Ethics statement
+The published data were collected and distributed in Japan and therefore comply with Japanese law.
+The legality of the collection and distribution process of J-CHAT was reviewed by a third-party Japanese lawyer.
+Under Article 30-4 of the Japanese Copyright Act, copyrighted materials may be used without explicit permission when the purpose is limited to “information analysis.”2 Accordingly, the dataset is released strictly for purposes related to information analysis, and not for entertainment or content redistribution.
+To mitigate privacy concerns, we performed anonymization of source metadata.
+In addition, the original recordings were segmented into dialogue units, preventing reconstruction of the original content for entertainment purposes.
+We also implement an opt-out policy, following prior large-scale web corpus practices (e.g.
+Common Crawl), allowing individuals or rights holders to request removal of their data from the dataset.
+8.  Acknowledgements
 This work was supported by AIST KAKUSEI project (FY2023), JST Moonshot JPMJMS2011 and JST FOREST JPMJFR226V.
-Appendix A dGSLM Training Details
-For the implementation of dGSLM, we used fairseq [^3] https://github.com/facebookresearch/fairseq repository by Meta which includes official implementation of the model.
-This repository is licensed as MIT.
-The number of train/valid/test is 1,003,397/9,991/100 dialogues for YouTube subset and for podcast subset it is 3,885,007/38,902/100 dialogues.
-For HuBERT model used in the speech-to-unit, we used the pretrained Japanese HuBERT model Sawada et al.
-(2024) [^4] https://huggingface.co/rinna/japanese-hubert-base.
-For k-means clustering, we used the original implementation of the dGSLM except for the number of clusters which is set to 1,000.
-The dGSLM model was configured as in the previous research Nguyen et al. (2023). The number of model parameters in each dGSLM model is 23,638,529.
-For training dGSLM model, we used 32 NVIDIA V100 GPUs with the learning rate of $2\times 10^{-4}$.
-The training was performed for 100,000 steps.
-The total duration of training was around 48 hours.
-The batch size was 36,864 tokens and the the maximum number of tokens per sample are set to 3,000.
-This is equivalent to the 1 minute of dialogue.
-For other hyper-parameters regarding the training, we followed the original paper Nguyen et al.
-(2023).
-For the XVector conditioned vocoder, we used the pretrained XVector model [^5] https://huggingface.co/speechbrain/spkrec-xvect-voxceleb.
-The vocoder was trained with the JVS Takamichi et al.
-(2020) and JVNV Xin et al.
-(2024a) corpus.
-These two corpora include Japanese reading-style, studio-quality speech without/with non-verbal expression, respectively.
-Training of vocoder took 2 days with 4 NVIDIA V100 GPUs.
-Appendix B Subjective Evaluation of the dGSLM Outputs
-B.1 Participant Details
-Participants of the each subjective evaluation were recruited on https://lancers.jp.
-They are paid with 75 Japanese yen upon the completion of the test.
-Given that the minimum wage of Japan is 1,004 Japanese yen and the test took approximately 4 minutes, the payment is adequate.
-Before the participation, the participants were instructed with the research purpose and how the collected data will be used in this research.
-The data collection protocol is approved by our institution’s ethics review board.
-Figure 3: Screenshot of instruction given to the participants on the naturalness MOS evaluation.
-Figure 4: Screenshot of instruction given to the participants on the meaningfulness MOS evaluation.
-Figures  3 and 4 show the screenshots of the instructions given to the participants for the naturalness and meaningfulness MOS evaluations, respectively.
-The English translation of the instruction for the naturalness MOS evaluation is as follows:
-Listen to the audio and rate the naturalness of the sound on a scale of 1 to 5 (very unnatural (1) to very natural (5)).
-Please conduct the evaluation using headphones in a quiet environment.
-Please note that some audio samples may have degraded sound quality, so please be careful with the volume level.
-Do not use your browser’s “back” or “reload” buttons during the test.
-The English translation of the instruction for the meaningfulness MOS evaluation is as follows:
-Listen to the audio and rate the meaningfulness of the dialogue on a 5-point scale from very meaningless (1) to very meaningful (5).
-Please evaluate in a quiet environment using headphones.
-Please note that some of the audio samples may have degraded sound quality, so please be careful with the volume.
-Please do not use your browser’s “back” or “reload” buttons during the test.
-B.2 P-values from the Subjective Evaluation
-Figure 5: P-values from the subjective evaluation on the naturalness of the dGSLM output samples
-Figure 6: P-values from the subjective evaluation on the meaningfulness of the dGSLM output samples
-Figure 5 and Figure 6 shows the p-values from the subjective evaluation on the dGSLM output conducted on Section 5.2. Specifically, we performed two sided T-test on the MOS score.
+9.  Bibliographical References
+  • V.
+Agrawal, A.
+Akinyemi, K.
+Alvero, M.
+Behrooz, J.
+Buffalini, F.
+M.
+Carlucci, J.
+Chen, J.
+Chen, Z.
+Chen, S.
+Cheng, et al.
+(2025) Seamless interaction: dyadic audiovisual motion modeling and large-scale dataset.
+arXiv preprint arXiv:2506.22554.
+Cited by: §1, §1, §2.
+  • C. Alexandra and G. Zipperlen (1996) CALLHOME Japanese speech LDC96S37. In Linguistic Data Consortium, Cited by: §1, Figure 2, §4.3.
+  • Z. Borsos, M. Sharifi, D. Vincent, E. Kharitonov, N. Zeghidour, and M. Tagliasacchi (2023) SoundStorm: efficient parallel audio generation. External Links: 2305.09636 Cited by: §1.
+  • G.
+Chen, S.
+Chai, G.
+Wang, J.
+Du, W.
+Zhang, C.
+Weng, D.
+Su, D.
+Povey, J.
+Trmal, J.
+Zhang, M.
+Jin, S.
+Khudanpur, S.
+Watanabe, S.
+Zhao, W.
+Zou, X.
+Li, X.
+Yao, Y.
+Wang, Z.
+You, and Z.
+Yan (2021) GigaSpeech: An Evolving, Multi-Domain ASR Corpus with 10,000 Hours of Transcribed Audio.
+In Proc.
+Interspeech 2021, pp.
+3670–3674.
+External Links: Document, ISSN 2958-1796 Cited by: §1, §1, §2.
+  • C.
+Cieri, D.
+Miller, and K.
+Walker (2004) The Fisher Corpus: a resource for the next generations of speech-to-text.
+In Proceedings of the Fourth International Conference on Language Resources and Evaluation (LREC’04), M.
+T.
+Lino, M.
+F.
+Xavier, F.
+Ferreira, R.
+Costa, and R.
+Silva (Eds.), Lisbon, Portugal.
+External Links: Link Cited by: §1, §2.
+  • A.
+Défossez, L.
+Mazaré, M.
+Orsini, A.
+Royer, P.
+Pérez, H.
+Jégou, E.
+Grave, and N.
+Zeghidour (2024) Moshi: a speech-text foundation model for real-time dialogue.
+arXiv arXiv:2410.00037.
+Cited by: §3.4.
+  • W.
+Hsu, B.
+Bolte, Y.
+H.
+Tsai, K.
+Lakhotia, R.
+Salakhutdinov, and A.
+Mohamed (2021) HuBERT: self-supervised speech representation learning by masked prediction of hidden units.
+IEEE/ACM Transactions on Audio, Speech, and Language Processing 29 (), pp.
+3451–3460.
+External Links: Document Cited by: Figure 3, §4.3.
+  • R.
+Huang, M.
+Li, D.
+Yang, J.
+Shi, X.
+Chang, Z.
+Ye, Y.
+Wu, Z.
+Hong, J.
+Huang, J.
+Liu, et al.
+(2024) Audiogpt: understanding and generating speech, music, sound, and talking head.
+In Proceedings of the AAAI Conference on Artificial Intelligence, Vol.
+38, pp.
+23802–23804.
+Cited by: §1.
+  • E.
+Kharitonov, A.
+Lee, A.
+Polyak, Y.
+Adi, J.
+Copet, K.
+Lakhotia, T.
+A.
+Nguyen, M.
+Riviere, A.
+Mohamed, E.
+Dupoux, and W.
+Hsu (2022) Text-free prosody-aware generative spoken language modeling.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), S.
+Muresan, P.
+Nakov, and A.
+Villavicencio (Eds.), Dublin, Ireland, pp.
+8666–8681.
+External Links: Link, Document Cited by: §5.1.
+  • Y.
+Koizumi, H.
+Zen, S.
+Karita, Y.
+Ding, K.
+Yatabe, N.
+Morioka, M.
+Bacchiani, Y.
+Zhang, W.
+Han, and A.
+Bapna (2023) LibriTTS-R: a restored multi-speaker text-to-speech corpus.
+In Proc.
+INTERSPEECH 2023, pp.
+5496–5500.
+External Links: Document, ISSN 2958-1796 Cited by: §3.3.
+  • J.
+Kong, J.
+Kim, and J.
+Bae (2020) HiFi-GAN: generative adversarial networks for efficient and high fidelity speech synthesis.
+In Advances in Neural Information Processing Systems, H.
+Larochelle, M.
+Ranzato, R.
+Hadsell, M.F.
+Balcan, and H.
+Lin (Eds.), Vol.
+33, pp.
+17022–17033.
+External Links: Link Cited by: §5.1.
+  • K.
+Lakhotia, E.
+Kharitonov, W.
+Hsu, Y.
+Adi, A.
+Polyak, B.
+Bolte, T.
+Nguyen, J.
+Copet, A.
+Baevski, A.
+Mohamed, et al.
+(2021) On generative spoken language modeling from raw audio.
+Transactions of the Association for Computational Linguistics 9, pp.
+1336–1354.
+External Links: Link, Document Cited by: §5.
+  • K.
+Lee, K.
+Park, and D.
+Kim (2023) DailyTalk: spoken dialogue dataset for conversational text-to-speech.
+In ICASSP 2023 - 2023 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), Vol.
+, pp.
+1–5.
+External Links: Document Cited by: §1.
+  • K. Mitsui, Y. Hono, and K. Sawada (2023) Towards human-like spoken dialogue generation between ai agents from written dialogue. External Links: 2310.01088 Cited by: §1.
+  • G.
+Mittag, B.
+Naderi, A.
+Chehadi, and S.
+Möller (2021) NISQA: A Deep CNN-Self-Attention Model for Multidimensional Speech Quality Prediction with Crowdsourced Datasets.
+In Proc.
+Interspeech 2021, pp.
+2127–2131.
+External Links: Document, ISSN 2958-1796 Cited by: §4.2.
+  • T.
+A.
+Nguyen, E.
+Kharitonov, J.
+Copet, Y.
+Adi, W.
+Hsu, A.
+Elkahky, P.
+Tomasello, R.
+Algayres, B.
+Sagot, A.
+Mohamed, and E.
+Dupoux (2023) Generative spoken dialogue language modeling.
+Transactions of the Association for Computational Linguistics 11, pp.
+250–266.
+External Links: Link, Document Cited by: §1, §1, §5.2, §5.2, §5.3, §5.
+  • S.
+Park, C.
+Kim, H.
+Rha, M.
+Kim, J.
+Hong, J.
+Yeo, and Y.
+Ro (2024) Let’s go real talk: spoken dialogue model for face-to-face conversation.
+In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), L.
+Ku, A.
+Martins, and V.
+Srikumar (Eds.), pp.
+16334–16348.
+External Links: Link, Document Cited by: §1, §2.
+  • A.
+Plaquet and H.
+Bredin (2023) Powerset multi-class cross entropy loss for neural speaker diarization.
+In Proc.
+INTERSPEECH 2023, pp.
+3222–3226.
+External Links: Document, ISSN 2958-1796 Cited by: §3.2.2.
+  • S.
+Rouard, F.
+Massa, and A.
+Défossez (2023) Hybrid transformers for music source separation.
+In ICASSP 2023 - 2023 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), Vol.
+, pp.
+1–5.
+External Links: Document Cited by: §3.3.
+  • Y.
+Saito, Y.
+Nishimura, S.
+Takamichi, K.
+Tachibana, and H.
+Saruwatari (2022) STUDIES: Corpus of Japanese Empathetic Dialogue Speech Towards Friendly Voice Agent.
+In Proc.
+Interspeech 2022, pp.
+5155–5159.
+External Links: Document, ISSN 2958-1796 Cited by: §1, §2, Figure 2, §4.3.
+  • K.
+Sawada, T.
+Zhao, M.
+Shing, K.
+Mitsui, A.
+Kaga, Y.
+Hono, T.
+Wakatsuki, and K.
+Mitsuda (2024) Release of pre-trained models for the Japanese language.
+In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), N.
+Calzolari, M.
+Kan, V.
+Hoste, A.
+Lenci, S.
+Sakti, and N.
+Xue (Eds.), Torino, Italia, pp.
+13898–13905.
+External Links: Link Cited by: §5.2.
+  • K.
+Seki, S.
+Takamichi, T.
+Saeki, and H.
+Saruwatari (2023) TEXT-to-speech synthesis from dark data with evaluation-in-the-loop data selection.
+In ICASSP 2023 - 2023 IEEE International Conference on Acoustics, Speech and Signal Processing, Vol.
+, pp.
+1–5.
+External Links: Document Cited by: §1.
+  • Z.
+Sheikh, S.
+Shimizu, S.
+Arora, J.
+Shi, S.
+Cornell, X.
+Li, and S.
+Watanabe (2025) Scalable spontaneous speech dataset (SSSD): crowdsourcing data collection to promote dialogue research.
+In Proc.
+Interspeech 2025, pp.
+3963–3967.
+Cited by: §1, §2.
+  • D.
+Snyder, D.
+Garcia-Romero, G.
+Sell, D.
+Povey, and S.
+Khudanpur (2018) X-Vectors: robust dnn embeddings for speaker recognition.
+In 2018 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), Vol.
+, pp.
+5329–5333.
+External Links: Document Cited by: §5.1.
+  • S.
+Takamichi, L.
+Kürzinger, T.
+Saeki, S.
+Shiota, and S.
+Watanabe (2021) JTubeSpeech: corpus of japanese speech collected from youtube for speech recognition and speaker verification.
+External Links: 2112.09323 Cited by: §1, §3.1.
+  • S.
+Takamichi, R.
+Sonobe, K.
+Mitsui, Y.
+Saito, T.
+Koriyama, N.
+Tanji, and H.
+Saruwatari (2020) JSUT and JVS: free japanese voice corpora for accelerating speech synthesis research.
+Acoustical Science and Technology 41 (5), pp.
+761–768.
+External Links: Document Cited by: §5.2, §5.3.
+  • D.
+Xin, J.
+Jiang, S.
+Takamichi, Y.
+Saito, A.
+Aizawa, and H.
+Saruwatari (2024) JVNV: a corpus of japanese emotional speech with verbal content and nonverbal expressions.
+IEEE Access 12 (), pp.
+19752–19764.
+External Links: Document Cited by: §5.2.
+  • H.
+Zen, V.
+Dang, R.
+Clark, Y.
+Zhang, R.
+J.
+Weiss, Y.
+Jia, Z.
+Chen, and Y.
+Wu (2019) LibriTTS: a corpus derived from LibriSpeech for text-to-speech.
+arXiv preprint arXiv:1904.02882.
+Cited by: §1, §2.